OpenStudy (anonymous):

Find the exact value by using a half-angle identity. sin(5π/12)

OpenStudy (anonymous):

\[|\sin \theta/2| = \sqrt{\frac{ 1-\cos \theta }{ 2 }}\] Let 5pi/12 be your theta, so plug in: \[| \sin 5\pi / 12 | = \sqrt{ .5 (1- \cos 5\pi / 6)}\] Can you solve it from here?